# A kinematic origin of the white dwarfs in the Solar neighborhood

### v1

@@ 5.2.3 Selection criteria effects
We have stated that our Gaia 100 pc sample can be considered effectively complete. However, it is well known that the overall completeness of the sample rapidly decreases for red colors and faint magnitudes (see, for intance, fig. 5 from Jiménez-Esteban et al., 2018; Gentile Fusillo et al., 2021). It is to be expected that this bias will have an effect on the average values of the observed velocities, as it is more likely that older objects will be excluded from the observed sample.
In order to check the effect of the selection criteria imposed by Gaia, we associate to each synthetic star a mass and a total age according to the Galactic models described in Torres et al. (2019b). By using appropriate cooling sequences from the set of evolutionary models of the La Plata group (Althaus et al., 2013; Camisassa et al., 2016, 2019), we can assign Gaia magnitudes to each simulated white dwarf. Furthermore, our Monte Carlo simulator is able to emulate the Gaia performance. We finally introduce a selection cut for those objects with astrometric errors and photometric flux errors larger than 10% and an apparent magnitude in Gaia G filter below 21 mag. Thus, the synthetic objects in our initial sample are split into those that pass Gaia’s selection criteria, and thus would be observable, and those that are excluded.
The results listed in Table 7 show that, whereas the majority of the thin disk objects can be observed by Gaia, a much larger fraction of halo objects can not. Thick disk objects are just in the middle, with about 50% of them not being observable by Gaia.
@@ 6 Conclusions
We have studied the kinematics and the orbital parameters of a synthetic white dwarf population, which was modeled from the observed properties of local white dwarfs within 100 pc away from the Sun. The simulated population was divided into thin and thick disk, and halo components, in order to compare with the results of the Random Forest classification of the Gaia sample (Torres et al., 2019b).
@@
By numerically integrating the Galactic orbits of simulated and observed stars, we have showed that the Solar neighborhood is a temporary concept, with an average lifetime of 0.6–3.3 Myr, depending on the average speed of the considered population. Nevertheless, most of local white dwarfs that belong to the thin disk, seem to originate from a relatively small annulus of about kpc in width that is centered on the LSR; thus, implying that most of thin disk white dwarfs have chemically and evolutionarily similar progenitors. The analysis of the average orbits and the distributions of guiding-center radii for thick disk and halo stars is characterized by a much broader coverage of the available space parameters, relating that to the radial mixing of stars. It is well established that these effects are likely due to the interaction with the Galactic bar and/or spiral arms potentials. We note a significant correlation among eccentric/retrograde and less-eccentric/prograde orbits of halo stars and their distribution of guiding-center radii, which possibly suggests a different origin either from accreted satellite galaxies or the existence of an in-situ stellar population in the inner halo.
@@
By simulating possible detection biases, which are either due to the Gaia detection sensitivity or to the kinematic classification of the observed sample, we showed that their importance is relatively small. In fact, the good agreement between simulated and observed crossing times of the Solar neighborhood are indicative that there is small contamination among kinematic populations. The largest divergence between lifetimes for the observed and simulated samples is noted for the thick disk population, for which we do not model the more complex kinematic signature of stellar streams. Moreover, the Gaia detection bias does not affect the overall parameter space, which allowed us a reliable coverage of the simulated distributions.
@@
The bias introduced by the known lack of accurate radial velocities for white dwarfs was previously discussed by Torres et al. (2019b) and Rowell & Kilic (2019), could account for systematic differences that could be rather large for thick disk and halo stars. However, the comparison with local main-sequence and sub-giant stars suggests that the systematic uncertainties affecting the estimates of the Solar neighborhood lifetime and space volume occupied by the Galactic orbits of the local sample should be in the range of 15–30 %. Nevertheless, we stress the importance of measuring accurate radial velocities of white dwarfs and we expect that such an issue will be partly resolved in the next 5–10 years thanks to the spectroscopic observations of the majority of Gaia white dwarfs by the WEAVE (Jin et al., 2024) and 4MOST (de Jong et al., 2012) surveys, which will additionally provide metal abundances for a sub-sample of white dwarfs in wide binary systems with a non-degenerate companion (Toloza et al., 2023). Such larger datasets will enable a more detailed analysis of the existing correlations among age, kinematics, and metallicity, which we recently studied for magnitude-limited samples of white dwarfs (e.g., Rebassa-Mansergas et al., 2021; Raddi et al., 2022), and are known to be related to the dynamical evolution of our own Galaxy.
@@ Acknowledgements.
- Garcia-Berro et al. (1998) Garcia-Berro, E., Torres, S., Isern, J., & Burkert, A. 1998, arXiv e-prints, astro

### v2

@@ 5.2.2 Effect of null radial velocity assumption
Note: A comparison with main sequence (MS) and sub- | ||||
giant (SG) stars within 100 pc from the Sun is included. |
@@ 5.2.3 Selection criteria effects
We have stated that our Gaia 100 pc sample can be considered effectively complete. However, it is well known that the overall completeness of the sample rapidly decreases for red colors and faint magnitudes (see, e.g., Fig. 5 from Jiménez-Esteban et al., 2018; Gentile Fusillo et al., 2021). It is to be expected that this bias will have an effect on the average values of the observed velocities, as it is more likely that older objects will be excluded from the observed sample.
To check the effect of the selection criteria imposed by Gaia, we associate to each synthetic star a mass and a total age according to the Galactic models described in Torres et al. (2019b). By using appropriate cooling sequences from the set of evolutionary models of the La Plata group (Althaus et al., 2013; Camisassa et al., 2016, 2019), we can assign Gaia magnitudes to each simulated white dwarf. Furthermore, our MC simulator is able to emulate the Gaia performance. Finally, we introduce a selection cut for those objects with astrometric errors and photometric flux errors larger than 10% and an apparent magnitude in Gaia G filter below 21 mag. Thus, the synthetic objects in our initial sample are split into those that pass Gaia’s selection criteria, and thus would be observable, and those that are excluded.
The results listed in Table 7 show that while the majority of the thin-disk objects can be observed by Gaia, a much larger fraction of halo objects can not. Thick disk objects are just in the middle, with about 50% of them not being observable by Gaia.
@@ 6 Conclusions
We have studied the kinematics and the orbital parameters of a synthetic white dwarf population, which was modeled from the observed properties of local white dwarfs within 100 pc away from the Sun. The simulated population was divided into thin- and thick-disk, and halo components to compare with the results of the Random Forest classification of the Gaia sample (Torres et al., 2019b).
@@
By numerically integrating the Galactic orbits of simulated and observed stars, we have showed that the solar neighborhood is a temporary concept, with an average lifetime of 0.6–3.3 Myr, depending on the average speed of the considered population. Nevertheless, most of the local white dwarfs that belong to the thin disk seem to originate from a relatively small annulus of about kpc in width that is centered on the LSR; this implies that most of of the thin-disk white dwarfs have progenitors that are chemically and evolutionarily similar. The analysis of the average orbits and the distributions of guiding-center radii for thick disk and halo stars is characterized by a much broader coverage of the available space parameters, relating that to the radial mixing of stars. It is well established that these effects are likely due to the interaction with the Galactic bar and/or spiral arms potentials. We note a significant correlation among eccentric-and-retrograde and less-eccentric-and-prograde orbits of halo stars and their distribution of guiding-center radii. This possibly suggests a different origin either from accreted satellite galaxies or the existence of an in-situ stellar population in the inner halo.
@@
By simulating possible detection biases, which are either due to the Gaia detection sensitivity or to the kinematic classification of the observed sample, we show that their importance is relatively small. In fact, the good agreement between simulated and observed crossing times of the solar neighborhood are indicative that there is small contamination among kinematic populations. The largest divergence between lifetimes for the observed and simulated samples is noted for the thick disk population, for which we do not model the more complex kinematic signature of stellar streams. Moreover, the Gaia detection bias does not affect the overall parameter space, which allowed us a reliable coverage of the simulated distributions.
@@
The bias introduced by the known lack of accurate radial velocities for white dwarfs was previously discussed by Torres et al. (2019b) and Rowell & Kilic (2019), could account for systematic differences that could be rather large for thick disk and halo stars. However, the comparison with local main-sequence and sub-giant stars suggests that the systematic uncertainties affecting the estimates of the solar neighborhood lifetime and space volume occupied by the Galactic orbits of the local sample should be in the range of 15–30 %. Nevertheless, we stress the importance of measuring accurate radial velocities of white dwarfs. We expect that this issue will be partly resolved in the next five to ten years thanks to the spectroscopic observations of the majority of Gaia white dwarfs by the WEAVE (Jin et al., 2024) and 4MOST (de Jong et al., 2012) surveys, which will additionally provide metal abundances for a sub-sample of white dwarfs in wide binary systems with a non-degenerate companion (Toloza et al., 2023). Such larger datasets will enable a more detailed analysis of the existing correlations among age, kinematics, and metallicity, which we recently studied for magnitude-limited samples of white dwarfs (e.g., Rebassa-Mansergas et al., 2021; Raddi et al., 2022) and are known to be related to the dynamical evolution of our own Galaxy.
@@ Acknowledgements.
- García-Berro et al. (1999) García-Berro, E., Torres, S., Isern, J., & Burkert, A. 1999, MNRAS, 302, 173